OpenStudy (anonymous):

Find the circumference of the circle with the given radius or diameter. (Use 3.14 for π.) radius = 23 m a. 6,644.24 m b. 144.44 m c. 72.22 m d. 1,661.06 m

OpenStudy (anonymous):

C = pi * D C is circumference and D is diameter

OpenStudy (anonymous):

Diameter is 2 * the radius

OpenStudy (anonymous):

C=2pi*r C=2(3.14)(23) C=144.51 meters D=2*r D=2(23) D=46 meters :)